OpenStudy (anonymous):

Working together, two pumps can drain a certain pool in 4 hours. if it takes the older pump 12 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own? Do not do any rounding.

OpenStudy (anonymous):

This seems easy but it is mindfu*king me

OpenStudy (anonymous):

haha right?

OpenStudy (anonymous):

it's those steps, but i got it wrong when i tried it...

OpenStudy (anonymous):

@bbbbbbbbbbbbbbbbbb

OpenStudy (wolf1728):

Formula is: 4 Hours = Pump1 * Pump2 / (Pump1 + Pump2) 4 hours = 12 * pump2 / (12 + Pump2) 4 * Pump2 + 48 = 12*pump2 48 = 8 * pump2 pump2 = 6 hours Double Check Both working together (12 * 6) / (12 + 6) 72 / 18 = 4 hours

OpenStudy (anonymous):

Left on together, the cold and hot water faucets of a certain bathtub take 5 minutes to fill the tub. If it takes the cold water faucet 15 minutes to fill the tub by itself, how long will it take the hot water faucet to fill the tub on its own? Do not do any rounding. What about this one ? And thank you @wolf1728 !!

OpenStudy (wolf1728):

Both faucets = 5 minutes cold faucet = 15 minutes 5 = 15 * hot / (15 + hot) 75 + 5 * hot = 15 * hot 75 = 10 * hot hot faucet = 7.5 hours Double Check Both faucets (15*7.5) / (15+7.5) 112.5 / 22.5 = 5 minutes !! successful !!!

OpenStudy (wolf1728):

Gee? Any chance for a medal ???

OpenStudy (anonymous):

Sorry I had to step away. But much appreciated! thank you so much!! :D